# Can MLLMs Guide Me Home? A Benchmark Study on Fine-Grained Visual Reasoning from Transit Maps

### v2

@@ 6 Conclusion
We provide a comprehensive overview in the Appendix, covering key components of our dataset, methodology, and analysis. Specifically, we include the question templates used during dataset construction, sources of transit maps from $30$ cities, detailed descriptions of the evaluation algorithm, experimental setup, and in-depth analyses of both correct and erroneous cases.
In addition, we further discuss diverse image processing strategies employed by various MLLMs when handling high-resolution visual inputs. We also address the stated limitations of our work, potential broader impacts, ethical considerations related to the dataset, and licensing information for the MLLMs used in our experiments.
[…]
@@ A.1 Question Template Summary
Departure Stop: xx Station
Arrival Stop: xx Station
@@
Departure Stop: xx Station
Arrival Stop: xx Station
@@
Departure Stop: xx Station
Arrival Stop: xx Station
@@
Departure Stop: xx Station
Arrival Stop: xx Station
@@
Departure Stop: xx Station
Arrival Stop: xx Station
Via Stops: xx Station, xx Station, xx Station
@@
Departure Stop: xx Station
Arrival Stop: xx Station
Via Stops: xx Station
A.2 Map Source
@@
Table A1: Source links for city transit maps used in ReasonMap. We present a total of $30$ cities sourced from $13$ countries.
@@
Budapest Link Oslo Link Rome Link
Lisboa Link Geneva Link Dubai Link
Auckland Link Sydney Link Singapore Link
Kuala Lumpur Link Los Angeles Link Miami Link
New York Link Toronto Link Washington Link
Guiyang Link Shanghai Link Huhehaote (Hohhot) Link
Nanchang Link Nanning Link Shenzhen Link
Hangzhou Link Dalian Link Kunming Link
Hefei Link Beijing Link Changzhou Link
Jinan Link Xi’an Link Changshang Link
Appendix B Evaluation Details
@@ B.1 Correctness and Quality Evaluation
We present the detailed algorithms for evaluating answer correctness and quality in Section 4 (Algorithm 1 for correctness evaluation and Algorithm 2 for quality evaluation).
@@
       if route name not in the Map Metadata then
@@
             Calculate absolute difference ($error$) in the number of via stops;
@@
[…]
@@
        map_score $\leftarrow$ map_score $+$ Option(via_stop_score or num_via_stop_score);
/* $10$ for short question; $20$ for long question */
[…]
@@
[…]
@@
B.2 Details about Difficulty-Aware Weighting.
("easy", "easy"): $1.0$ ("middle", "easy"): $1.5$ ("hard", "easy"): $2.0$
("easy", "middle"): $1.5$ ("middle", "middle"): $2.0$ ("hard", "middle"): $2.5$
("easy", "hard"): $2.0$ ("middle", "hard"): $2.5$ ("hard", "hard"): $3.0$
Appendix C Case Analysis
We provide additional case analyses covering both correct and incorrect predictions, along with detailed comparisons of their respective reasoning processes.
We compare Doubao-415 and Doubao-428 (Figure A1), both of which reach the correct destination (from Augustins Station to Poterie Station) but via distinct reasoning paths. Doubao-415 correctly identifies early that both stations are on Line 18 and efficiently converges on the optimal, direct route without transfers. In contrast, Doubao-428 misclassifies Augustins as being on Line 12 and, assuming Poterie is on Line 18, proposes a transfer route via Plainpalais—functionally correct but suboptimal due to unnecessary complexity. Both models engage in extensive self-correction ($7270$ tokens for Doubao-428; $4474$ for Doubao-415), highlighting the significant downstream impact of early-stage misjudgments. Moreover, visual reasoning limitations persist: despite correctly recognizing Augustins on Line 12, Doubao-415 commits to a transfer path and fails to re-evaluate the possibility of a direct connection. This indicates room for improvement in both early visual grounding and global route optimality awareness.
We then analyze the observed pattern when comparing the full input and text-only variants in the case (in Figure A2). The model with full visual access accurately identifies both stations on the Yellow Line and outputs the optimal direct route with the correct number of via stops. In contrast, the text-only variant makes an early misclassification, placing both stations on the Blue Line (Azul) and constructing a plausible but entirely incorrect sequence of intermediate stops. Although the final answer format appears coherent, the underlying logic is flawed due to the initial error in line recognition. This further illustrates the importance of visual input in spatial reasoning tasks, where even minor misinterpretations can lead to fundamentally incorrect conclusions. Additionally, some models, such as the InternVL3 series, default to rejection when visual input is absent.
We further present several error cases (in Figure A3), where Doubao-415 still exhibits visual confusion. In contrast, Qwen2.5-VL-32B-I, when lacking visual input, behaves differently from the InternVL3 series: rather than rejecting the query outright, it attempts to reason over the available information without producing a final answer, while explicitly notifying the user of the missing visual input.
Appendix D Further Discussions
D.1 High-Resolution Image Preprocessing.
@@ Appendix D Further Discussions
Table A2: Comparison of high-resolution image preprocessing strategies across different MLLMs. We use “--” to denote unspecified or unclear content.
@@
Qwen2.5-VL series [2] 2D-RoPE ($2\times 2$ Concat + MLP)
QVQ-72B-Preview [20] 2D-RoPE ($2\times 2$ Concat + MLP)
InternVL3 series [3] V2PE (Unshuffle + MLP)
Kimi-VL series [16] 2D-RoPE (Shuffle + MLP)
Skywork-R1V-38B [17, 18] -
Gemini [50] (Tiling+Resize) -
Doubao-1.5-Pro series [19] 2D-RoPE ($2\times 2$ Pooling + MLP)
@@
D.2 Limitations and Future Work

### v3

@@ Acknowledgement
We provide a comprehensive overview in the Appendix, covering key details of our dataset, methodology, evaluation, training baseline, analysis, and further discussions. Specifically, we include the question templates, quality control details, a fine-grained taxonomy of difficulty, and sources of transit maps from $30$ cities for ReasonMap construction in Appendix A. We then report detailed descriptions of the evaluation algorithm, experimental setup, and GRPO training in Appendix B. In Appendix C, we include supplementary results and conduct more experiments, including supplementary results, evaluation of symbolic representation and an ablation study about languages. We also provide the results of fine-grained error analysis metrics and systematically analyze failure causes. In Appendix D, we further extend case analysis by providing more classical cases. In addition, we further discuss the stated limitations, future directions, and potential broader impacts of our work in Appendix E. We finally present public implementation for the MLLMs used in our experiments, LLM usage statement, and ethical statement (see Appendix F).
@@ A.1 Question Template Summary
Departure Stop: $x$ Station
Arrival Stop: $x$ Station
@@
Departure Stop: $x$ Station
Arrival Stop: $x$ Station
@@
Departure Stop: $x$ Station
Arrival Stop: $x$ Station
@@
Departure Stop: $x$ Station
Arrival Stop: $x$ Station
@@
Departure Stop: $x$ Station
Arrival Stop: $x$ Station
Via Stops: $x$ Station, $x$ Station
@@
Departure Stop: $x$ Station
Arrival Stop: $x$ Station
Via Stops: $x$ Station
A.2 A More Fine-grained Taxonomy of Difficulty
Beyond the easy, middle, and hard categorization for map and question difficulty, we provide three additional difficulty aware labels: 1) city_line_countcity\_line\_count, the total number of lines in a city (i.e., a proxy for map difficulty); 2) city_transfer_countcity\_transfer\_count, the total number of transfer stations in a city (i.e., a proxy for map difficulty); and 3) question_transfer_countquestion\_transfer\_count, the number of transfers in the queried route (i.e., a proxy for question difficulty). These labels enable fine-grained category design and filtering in subsequent analyses.
A.3 Quality Control Details
Our quality control combines automated checks with manual refinement. Specifically, we first validate route correctness (e.g., start stop, arrival stop, and connectivity), followed by manual checks to ensure visual consistency (i.e., routes can be inferred from maps). Questions or GTs with issues are corrected or removed. Three domain experts reviewed the data and identified an error rate of $\sim$1616%, after which all questions/GTs were corrected and verified to be accurate. Finally, we systematically adjust the difficulty distributions to prevent bias and ensure a balanced evaluation benchmark.
A.4 Map Source
@@
Table A1: Source links to the city transit maps used in the ReasonMap dataset. We present a total of $30$ cities sourced from $13$ countries.
@@
Budapest [Link] Oslo [Link] Rome [Link]
Lisboa [Link] Geneva [Link] Dubai [Link]
Auckland [Link] Sydney [Link] Singapore [Link]
Kuala Lumpur [Link] Los Angeles [Link] Miami [Link]
New York [Link] Toronto [Link] Washington [Link]
Guiyang [Link] Shanghai [Link] Huhehaote (Hohhot) [Link]
Nanchang [Link] Nanning [Link] Shenzhen [Link]
Hangzhou [Link] Dalian [Link] Kunming [Link]
Hefei [Link] Beijing [Link] Changzhou [Link]
Jinan [Link] Xi’an [Link] Changshang [Link]
Appendix B Details of Evaluation and Training Baseline
@@ B.1 Correctness and Quality Evaluation
We present the detailed algorithms for evaluating answer correctness and quality (Algorithm 1 for correctness evaluation and Algorithm 2 for quality evaluation).
For matching (e.g., ==) in the algorithms, we apply rule-based corrections on top of string matching to account for semantically irrelevant formatting variations (e.g., Line 1” == Route 1” == “1”), preventing evaluation failures caused solely by stylistic or linguistic differences. These corrections are deliberately limited to remain consistent with the format requirements of each question.
Additionally, for multilingual maps, the pipeline is identical to that of English maps, except that route and station names are retained in their local language. During evaluation, we accept both the local language and its English translation as correct, prioritizing semantic correctness.
@@
  if route name not in the Metro Data then
@@
     Calculate absolute difference (errorerror) in the number of via stops;
@@
  via_stop_score $\leftarrow\min(10,\texttt{via\_stop\_score})$;
  num_via_stop_score $\leftarrow\min(10,\texttt{num\_via\_stop\_score})$;
@@
   map_score $\leftarrow$ map_score ++ Option(via_stop_score or num_via_stop_score);
/* $10$ for short question; 2020 for long question */
map_score $\leftarrow\min(10,\texttt{map\_score})/min(20,\texttt{map\_score})$ ;
@@
  map_score $\leftarrow\texttt{map\_score}+10/\texttt{map\_score}+20$;
@@
B.2 High-Resolution Image Preprocessing.
@@
Table A2: Comparison of high-resolution image preprocessing strategies across different MLLMs. We use “-” to denote unspecified or unclear content.
@@
    Qwen2.5-VL series [3]     ✓     2D-RoPE     ✓ ($2\times 2$ Concat + MLP)
    QVQ-72B-Preview [51]     ✓     2D-RoPE     ✓ ($2\times 2$ Concat + MLP)
    InternVL3 series [102]     ✓     V2PE     ✓ (Unshuffle + MLP)
    Kimi-VL series [64]     ✓     2D-RoPE     ✓ (Shuffle + MLP)
    Skywork-R1V-38B [74, 49]     ✓     -     ✗
    Gemini [19]     ✗ (Tiling+Resize)     -     ✗
    Doubao-1.5-Pro series [6]     ✓     2D-RoPE     ✓ ($2\times 2$ Pooling + MLP)
@@
B.3 Details about Difficulty-Aware Weighting.
Difficulty Pair Weight
(“easy”, “easy”) 1.0
(“medium”, “easy”) 1.5
(“hard”, “easy”) 2.0
(“easy”, “medium”) 1.5
(“medium”, “medium”) 2.0
(“hard”, “medium”) 2.5
(“easy”, “hard”) 2.0
(“medium”, “hard”) 2.5
(“hard”, “hard”) 3.0
B.4 Details of GRPO RL Training
GRPO [57] extends standard policy gradient methods by normalizing rewards within a sampled group, which stabilizes optimization and encourages relative preference learning. Specifically, given an input $x$ and a group of $K$ sampled outputs $G=\{y_{i}\}_{i=1}^{K}$ with their corresponding scalar rewards $\{r_{i}\}_{i=1}^{K}$, the centered group advantage $\hat{A}_{i}$ is computed as the deviation of each sample’s reward from the group mean:
$\hat{A}_{i}=r_{i}-\frac{1}{K}\sum_{j=1}^{K}r_{j}.$ (1)
The policy parameters $\theta$ are then updated to maximize the following objective:
$\max_{\theta}\;\mathcal{L}(\theta)=\sum_{i=1}^{K}\hat{A}_{i}\log\pi_{\theta}(y_{i}\mid x),$ (2)
where $\pi_{\theta}(y_{i}\mid x)$ denotes the model likelihood of generating $y_{i}$ under parameters $\theta$. This objective encourages the model to increase the probability of outputs with above-average rewards while suppressing those with below-average ones. In our implementation, the reward $r_{i}$ is composed of an accuracy component and a format component.
Appendix C Supplementary Experiments
C.1 Supplementary Results
We provide additional visualization results in Figure A1 and A2 to better illustrate our evaluation on ReasonMap as follows. Figure A1 and A2 illustrate the model’s accuracy across different difficulty levels and cities, respectively. As shown, accuracy decreases with increasing difficulty and varies considerably across cities.
Refer to caption Refer to caption
(a) Accuracy for short questions (b) Accuracy for long questions
Figure A1: Accuracy across difficulty combinations for four representative MLLMs (Qwen2.5-VL-72B-I, InternVL3-78B, OpenAI o3, and Doubao-415). Each difficulty combination is denoted by a pair (e.g., easy-hard), where the first term indicates question difficulty and the second term represents map difficulty. The pair (hard-middle) contains only one sample, leading to an accuracy of 100%. We summarize the number of evaluation samples in each difficulty bucket: $55$ samples for easy-easy, $46$ for easy-middle, $28$ for middle-easy, 77 for hard-easy, $23$ for middle-middle, $80$ for easy-hard, 11 for hard-middle, $57$ for middle-hard, and $15$ for hard-hard.
Refer to caption Refer to caption
(a) Accuracy for short questions (b) Accuracy for long questions
Figure A2: Accuracy across different cities for four representative MLLMs (Qwen2.5-VL-72B-I, InternVL3-78B, OpenAI o3, and Doubao-415). Each city is marked with the corresponding map difficulty and the country flag. Each city in the test set provides a specific number of samples per model: $32$ samples for Auckland, $34$ for Los Angeles, 77 for Miami, $35$ for Lisboa, $18$ for Geneva, $40$ for Beijing, $39$ for Hangzhou, $17$ for Budapest, $39$ for Singapore, $40$ for Rome, and $11$ for Toronto.
C.2 Fine-grained Error Analysis Metric Summary
We report multiple fine-grained error analysis metrics in Table A3 as follows: (1) deparrscoredep-arr~score: +1+1 if both the start and end stations are correct; (2) routenamescoreroute~name~score: +2+2 for each correctly identified line name along the route; (3) stopsscorestops~score: +1+1 for each correctly identified intermediate stop; (4) num_via_stop_scorenum\_via\_stop\_score (only for long questions): computed by taking the absolute difference between the number of via stops in the answer and the reference route, and mapping it to a score from 0 to $4$; (5) via_stop_scorevia\_stop\_score (only for long questions): calculated by averaging the number of correctly matched via stops (up to $10$) and the Intersection-over-Union (IoU) between the via stop sets of the answer and reference route (scaled to $10$).
Table A3: Fine-grained error analysis metrics of various MLLMs. $S.$ represents results for short questions, while $L.$ denotes results for long questions. Bold indicates the best results among open-source and closed-source models, respectively.
Model Type Dep-Arr Score ($S.$ / $L.$) Route Name Score ($S.$ / $L.$) Stops Score ($S.$ / $L.$) Num. Via Stop Score ($L.$) Via Stop Score ($L.$)
Open-source Models
Qwen2.5-VL-3B-Instruct [3] Base 0.86 / 0.78 0.03 / 0.02 1.03 / 0.96 0.42 0.00
Qwen2.5-VL-32B-Instruct [3] Base 0.95 / 0.92 0.09 / 0.10 1.16 / 1.19 1.57 0.01
Qwen2.5-VL-72B-Instruct [3] Base 0.96 / 0.95 0.22 / 0.24 1.23 / 1.22 1.56 0.04
InternVL3-38B [102] Base 0.87 / 0.84 0.06 / 0.10 1.08 / 1.12 1.63 0.00
InternVL3-78B [102] Base 0.96 / 0.89 0.15 / 0.17 1.15 / 1.12 1.46 0.01
Kimi-VL-A3B-Instruct [64] Base 0.89 / 0.88 0.07 / 0.07 1.06 / 1.11 0.91 0.02
Kimi-VL-A3B-Thinking [64] Reasoning 0.80 / 0.65 0.08 / 0.10 0.99 / 0.79 0.50 0.00
Skywork-R1V-38B [74] Reasoning 0.60 / 0.62 0.06 / 0.09 0.74 / 0.71 1.00 0.00
QvQ-72B-Preview [51] Reasoning 0.35 / 0.22 0.03 / 0.02 0.42 / 0.29 0.20 0.01
Closed-source Models
Doubao-115 [6] Base 0.78 / 0.96 0.08 / 0.18 1.08 / 1.31 1.94 0.06
OpenAI 4o [46] Base 0.97 / 0.95 0.22 / 0.29 1.49 / 1.53 2.22 0.04
Doubao-415 [6] Reasoning 0.98 / 0.98 0.33 / 0.30 1.57 / 1.65 2.37 0.08
Doubao-428 [6] Reasoning 0.73 / 0.75 0.00 / 0.03 1.19 / 1.27 2.27 0.00
Gemini-2.5-Flash [19] Reasoning 0.93 / 0.67 0.27 / 0.29 1.67 / 1.22 1.82 0.05
OpenAI o3 [48] Reasoning 0.99 / 0.91 0.32 / 0.16 1.77 / 1.73 3.31 0.03
We further provide a sensitivity analysis with two additional weighting schemes for the components of the map score (C1 & C2 in Table A4). Based on the results in Tab. A3A4, performance ranking remains consistent across different weighting schemes, while all components of the map score exhibit similar increasing or decreasing trends.
Table A4: Ablation on the map score. MS denotes the map score. C1 uses the avg scheme, while C2 excludes the Dep-Arr score part.
Model Weighted MS (S./L.) C1 MS (S./L.) C2 MS (S./L.)
Qwen2.5-VL-32B-I $3.88$ / $6.84$ $3.85$ / $5.80$ $2.90$ / $4.88$
Qwen2.5-VL-72B-I $\mathbf{5.09}$ / $\mathbf{8.80}$ $\mathbf{5.08}$ / $\mathbf{8.83}$ $\mathbf{4.12}$ / $\mathbf{7.88}$
OpenAI GPT-4o $6.84$ / $13.57$ $6.80$ / $13.59$ $5.82$ / $12.64$
OpenAI o3 $\mathbf{9.53}$ / $\mathbf{17.96}$ $\mathbf{9.38}$ / $\mathbf{17.95}$ $\mathbf{8.39}$ / $\mathbf{17.04}$
C.3 Further Experiments about Languages
We conduct an ablation study under the textualized representation paradigm (as mentioned in Appendix C.4). In this setting, visual images are not involved, which allows us to safely replace all non-English station names with unique English aliases without introducing visual inconsistencies. This approach isolates the language prior factor and avoids any potential confounding effects from visual modifications. Concretely, we manually replace all Chinese station names in Beijing and Hangzhou with unique English station names (e.g., mapping them to New York stops: ‘zhichunli’ <<->> 86 St), preserving the original transit map structure. The results under this setting are as follows.
Table A5: Evaluations on Beijing and Hangzhou (with and without English). $S.$ represents results for short questions, while $L.$ denotes results for long questions. Bold indicates performance improvements, while italicized values represent performance degradation.
Model Beijing (S. / L.) Beijing (w. English) (S. / L.) Hangzhou (S. / L.) Hangzhou (w. English) (S. / L.)
Kimi-VL-A3B-Instruct [64] 36.76% / 17.30% 23.78% / 20.81% 40.00% / 42.22% 42.22% / 45.95%
Doubao-115 [21] 64.86% / 50.51% 45.95% / 52.70% 82.22% / 64.44% 67.78% / 65.56%
Doubao-415 [21] 84.86% / 74.05% 88.65% / 85.95% 94.44% / 97.22% 87.78% / 100%
Overall, we observe from the results in Table A5 that using English labels leads to performance improvements, particularly for long-form questions. This suggests that the model indeed exhibits a language bias, with English showing an advantage over Chinese, which may be attributed to differences in pre-training data distributions.
C.4 Further Experiments about Symbolic Representation of Maps
We conduct further experiments about deterministic baselines derived from symbolic representations of the maps. This setting can serve as a theoretical performance ceiling, independent of perceptual challenges faced by MLLMs. We replace the visual input with symbolic representations extracted from the underlying map structure. Specifically, we convert all routes and station information into textual form to represent the topological structure of the map. This textualized representation is then used for evaluation. Specifically, we provide the model with textualized representations and the question as input, without including any visual maps.
Table A6: Evaluations of various MLLMs using symbolic representation. $S.$ represents results for short questions, while $L.$ denotes results for long questions. Bold indicates the best results among open-source and closed-source models, respectively.
Model Type Weighted Acc. (S. / L.) #Tokens (S. / L.)
Open-source Models
Qwen2.5-VL-3B-Instruct [3] Base 22.83% / 19.79% 51 / 162
Qwen2.5-VL-32B-Instruct [3] Base 25.52% / 18.77% 97 / 297
Kimi-VL-A3B-Instruct [64] Base 39.58% / 34.81% 43 / 55
Closed-source Models
Doubao-115 [6] Base 81.16% / 72.66% 41 / 82
OpenAI 4o [46] Base 82.38% / 78.91% 40 / 70
Doubao-415 [6] Reasoning 95.31% / 93.66% 563 / 1561
By comparing the results in Table A6 with those in Table 2 of the main paper, we observe a clear performance improvement. This is expected, as replacing the visual map with textualized representations substantially reduces task difficulty, as it removes the need to assess visual capabilities such as OCR and grounding. We further note that prior works, such as MapBench [77] and CityBench [13], also focus on visual map interpretation without constructing explicit symbolic baselines.
C.5 Further Systematic Analysis on Failure Causes
Case 1: grounding Q: Hollywood/Highland Station ->> Wilshire/Vermont Station? <<image:los-angeles.png>> A (Kimi-VL-A3B-Thinking): … Let me start by locating both stations on the map. Hollywood/Highland is on … A Line ()… Wilshire/Vermont is on the C Line () … () A (Doubao-415): … First, I need to see where these two sites are located on a map. Hollywood/Highland … on B Line () … Wilshire/Vermont … is on D Line and B Line () … Route Name: B Line Departure Stop: Hollywood/Highland Station Arrival Stop: Wilshire/Vermont Station ()
We systematically analyze failure causes, focusing on three MLLM capabilities pertinent to fine-grained visual reasoning (e.g., OCR, grounding, and spatial reasoning). To assess OCR capabilities, we collect metrics of 9 representative MLLMs on OCRBench [43]. Comparing these with their performance on ReasonMap as shown in Table A7 in the paper, we observe no clear correlation between OCR ability and ReasonMap accuracy. Notably, this trend holds across both open-source and closed-source models, suggesting that stronger OCR performance alone does not lead to better fine-grained visual reasoning. For instance, among open-source models, InternVL3-78B achieves the highest OCRBench scores, but underperforms Qwen2.5-VL-72B-Instruct on ReasonMap.
Case 2: spatial reasoning Q: Ponte Mammolo Station ->> Sta. Maria del Soccorso Station? <<image:rome.png>> A (Skywork-R1V-38B): … Line A: Orange Line B: Blue Line C: Green () … Ponte Mammolo Station … on Line B () … Let me trace the path of Line B from Ponte Mammolo towards the center of Rome () … But I need to go from Ponte Mammolo to Sta. Maria del Soccorso. Looking at the map, I don’t see a direct connection between these two stations via Line B. () So, I might need to transfer to another line … Let me check if there’s another line that connects to Sta. Maria del Soccorso … () A (Gemini-2.5-Flash): Route Name: Line B Departure Stop: Ponte Mammolo Station Arrival Stop: Sta. Maria del Soccorso Station Number of Via Stops: 0 ()
Table A7: Evaluations of various MLLMs on OCRBench. Bold indicates the best results among open-source and closed-source models, respectively. The references in the table indicate the result sources. All results are collected from the technical reports.
Model Type OCRBench
Open-source Models
Qwen2.5-VL-3B-Instruct [3] Base 797
Qwen2.5-VL-72B-Instruct [3] Base 885
InternVL3-38B [102] Base 886
InternVL3-78B [102] Base 906
Kimi-VL-A3B-Instruct [64] Base 864
Kimi-VL-A3B-Thinking [64] Reasoning 864
Closed-source Models
OpenAI 4o [64] Base 815
Doubao1.5-VL (non-thinking) [21] Base 881
Doubao1.5-VL (thinking) [21] Reasoning 861
We further conduct more in-depth case analyses, which reveal that the main causes of failure are grounding and spatial reasoning, as illustrated in the following example. We observe that OCR errors rarely occur, and most failure cases are instead caused by grounding or spatial reasoning issues.
For instance, in Case 1, Kimi-VL-A3B-Thinking incorrectly identifies the line of the departure station, indicating a grounding error that leads to subsequent reasoning failures. In Case 2, Skywork-R1V-38B correctly performs OCR and grounding in the initial steps, but fails in the reasoning stage (i.e., it does not prioritize locating the arrival station and instead attempts to construct incorrect indirect paths). Such failures reflect deficiencies in spatial reasoning, particularly in planning and executing core steps of pathfinding. These cases further indicate that the principal capability gap between open-source and closed-source models lies in grounding and spatial reasoning.
Appendix D Case Analysis
We provide additional case analyses covering both correct and incorrect predictions, along with detailed comparisons of their respective reasoning processes. We first compare Doubao-415 and Doubao-428 (Figure A3), both of which reach the correct destination (from Augustins Station to Poterie Station) but via distinct reasoning paths. Doubao-415 correctly identifies early that both stations are on Line 18 and efficiently converges on the optimal, direct route without transfers. In contrast, Doubao-428 misclassifies Augustins as being on Line 12 and, assuming Poterie is on Line 18, proposes a transfer route via Plainpalais—functionally correct but suboptimal due to unnecessary complexity. Both models engage in extensive self-correction , highlighting the significant downstream impact of early-stage misjudgments. Moreover, visual reasoning limitations persist: despite correctly recognizing Augustins on Line 12, Doubao-415 commits to a transfer path and fails to re-evaluate the possibility of a direct connection. This indicates room for improvement in both early visual grounding and global route optimality awareness. We then analyze the observed pattern when comparing the full input and text-only variants in the case (in Figure A4). The model with full visual access accurately identifies both stations on the Yellow Line and outputs the optimal direct route with the correct number of via stops. In contrast, the text-only variant makes an early misclassification, placing both stations on the Blue Line (Azul) and constructing a plausible but entirely incorrect sequence of intermediate stops. Although the final answer format appears coherent, the underlying logic is flawed due to the initial error in line recognition. This further illustrates the importance of visual input in spatial reasoning tasks, where even minor misinterpretations can lead to fundamentally incorrect conclusions. Additionally, some models, such as the InternVL3 series, default to rejection when visual input is absent. We further present several error cases in Figure A5, where Doubao-415 still exhibits visual confusion. In contrast, Qwen2.5-VL-32B-I, when lacking visual input, behaves differently from the InternVL3 series: rather than rejecting the query outright, it attempts to reason over the available information without producing a final answer, while explicitly notifying the missing visual input.
Appendix E Further Discussions
E.1 Limitations and Future Work